# Double-Signed Fragmented DNSSEC for Countering Quantum Threat

### v1

@@ 3.2. Quantum-Safe DNSSEC
Simply replacing the conventional digital signatures with post-quantum candidates mentioned above may not make the DNSSEC quantum-safe and may even make DNSSEC susceptible to attacks conducted using today’s classical computer (see the details mentioned above). Therefore, according to ENISA in (ENISA, 2022), a better and more cautious way to proceed is to use post-quantum digital signatures as an extra layer of protection in conjunction with conventional digital signatures. The combination essentially leads to a double-signed DNSSEC (note that, the terms double-signed, combiners, or hybrid cryptosystems are used interchangeably). A well-designed double-signed DNSSEC will offer the benefit that it will remain strong if at least one of the used digital signatures is secure against a specific attack. In a double-signed DNSSEC, we will accept the IP address in a resolution process only if both signatures are verified. In particular, any attack revealed on post-quantum digital signatures in the near future that can be conducted with today’s classic computer will almost always fail to comprise the address resolution process as the security of conventional digital signatures is established and assured against classical computers. Similarly, if the post-quantum digital signatures remain secure and no attacks can be conducted upon them, they will be secure against CRQC-driven attacks. Therefore, the double-signed approach is best suited for the interim period, as otherwise (i.e., with conventional or post-quantum digital signatures alone), DNSSEC may unknowingly be susceptible to attacks in the interim period as mentioned above.
@@
As far as the double signing is concerned, existing literature specifies two common interfaces (ENISA, 2022). The first one, referred to as detached-message interface simply concatenates a pre-quantum signature $X$ with a post-quantum signature $Y$. The verifier only accepts the results when the $X$ verifier accepts the $X$ signature and the $Y$ verifier accepts the $Y$ signature. The second interface, referred to as signed-message interface, typically signs with a pre-quantum algorithm to get the first signature $X$ and then signs the result - i.e., $X$ with the post-quantum algorithm to get the final signature (e.g, $Z$). The double-signature verifier in this interface normally applies the post-quantum verifier first followed by the pre-quantum verifier (or vice versa depending upon the signing agreement).
@@
For the double-signed DNSSEC, we in this paper adopt an approach akin to the detached-message interface. Our reason is that this approach is relatively easier than others to integrate into our BIND9-based setup (see details on evaluation setup in Section 5). The overview of our double-signed DNSSEC is shown in Figure. 7. It essentially integrates two signatures - i.e., pre-quantum and post-quantum in the signed zone files on the authoritative server. Both these signatures are verified independently on the resolver before accepting any received resource record as verified and trusted to offer enhanced security.
@@ 4. Double-Signatures and Maximum Message Size Issue
In this section, we discuss the message size issue when double signatures are incorporated in DNSSEC along with the fragmentation approach that we have devised and tested for accommodating the dual signatures in DNSSEC.
@@
Since DNS(SEC) is an application-layer protocol, the generated messages are first placed in UDP packets, which subsequently are placed in IP packets that make the payload of link layer frames. The size of these payloads has a limit that is dictated by the Maximum Transmission Unit (MTU) of the physical links that these frames traverse. The size of the frame’s payload greater than the link’s MTU generally enforces routers to fragment these frames into smaller IP packets that travel separately to the resolver. However, the underlying transport protocol - i.e., UDP is mainly unreliable for successful delivery of all these fragments and thus may lead to failure of address resolution. To circumvent this issue, either we need to resort to a reliable transport protocol - i.e., TCP or we need to formulate messages such that their size is less than the MTU of the majority of physical links over which these DNS(SEC) messages will likely travel. Since as per our discussion in Section 1, TCP has support issues in existing DNS infrastructure and also is not as lightweight as UDP (Van Den Broek et al., 2014; van Rijswijk-Deij et al., 2016; Mao et al., 2022), the only option is to restrict DNS message size such that IP-level fragmentation is not triggered. To accomplish this, initial RFCs on DNS (Mockapetris, 1987) specify the maximum DNS message size to be less than (or equal to) 512B. However, the Extension Mechanism for DNS - i.e., EDNS(0) as defined in RFC 6891 (Damas and Vixie, 2013), theoretically allows a maximum message size of up to 64KB. Abstractly, this mechanism introduces a pseudo-RR called OPTION (alluded as OPT) and allows resolvers to specify buffer size in the ‘Class’ field of this pseudo-RR in the ‘additional section’ of an RR (see RR format in Fig. 3). Therefore, through EDNS(0) it is theoretically possible to increase DNS message size to a maximum of XB if the MTU of the end-to-end path between resolver and name server is at least ‘X+8(UDP Header)+40(IPv6 Header). However, as of now, there exists no standard way to estimate the MTU of end-to-end links. Therefore, taking the cautious approach, the maximum DNS message size is limited to 1232B so as to avoid IP-level fragmentation and frequent TCP re-transmission (Manual, 2023) (assuming MTU of 1280B).
@@
The above-mentioned limit on the maximum DNS message size (i.e., 1232B) creates a particular challenge for double-signed DNSSEC as envisioned in this work. Precisely, any combination of pre-quantum and post-quantum signatures in double-signed DNSSEC will lead to response messages of size significantly larger than the limit discussed above. For example, our investigation shows that the message in response to a query of ‘Type A’ (see address resolution process depicted in Figure 7) is generally the smallest in size. However, our thorough analysis suggests that double signed DNSSEC with pre- and post-quantum algorithms that lead to the signatures of the smallest sizes (e.g., pre-quantum ECDSA of 64B and post-quantum FALCON512 of 690B), has an approximate response to this query of size 2500B (note that exact size depends upon variable fields in an RR). The other messages such as those that carry ‘DNSKEY RRs’ are even larger. Therefore, to accommodate double signatures in DNSSEC, we need to fragment these large responses at the application layer - i.e., within the DNS(SEC) protocol itself.
@@
There exist ways to fragment DNSSEC messages at the application layer in the post-quantum era. For example, ARRF (A Resource Record Fragment) mechanism (Goertzen and Stebila, 2022) proposes a request-based fragmentation approach at the application layer. For response messages larger than the threshold limit, ARRF on name servers ‘truncate’ the first fragments indicating to the resolver that there are additional fragments before this response can be re-assembled and considered complete. To facilitate the resolver in efficient reception of fragments, ARRF introduces a new pseudo-RR called Resource Record Fragments (RRFRAGS, akin to OPT described above). This pseudo-RR adopts the format of a typical RR as shown in Figure 3, however, its fields are customized such that the name server can indicate to the resolver in the first fragment how the complete response would look like once all the fragments are received. Likewise, it also facilitates name servers in ascertaining whether the resolver requests a fragment (or is it a normal query). Although ARRF seems to perform fragmentation reliably, it has two main shortcomings. The first one is that it relies on a non-standard RR that will potentially cause issues with the middleboxes. Additionally, as acknowledged by the authors of ARRF, this approach may also be susceptible to memory exhaustion attacks.
@@
Given the above-mentioned problems with ARRF, QName-Based Fragmentation (QBF) is proposed in (Rawat and Jhanwar, 2024). Like ARRF, QBF also adopts a request-based fragmentation approach but leverages standard RRs, unlike ARRF which relies on customized RRs for its operations. When the response size is bigger than the threshold, QBF sends a fraction of the signature or public key in the first fragment and sets the ‘TC’ flag in the header of the response to ‘1’, indicating to the resolver that this response is part of a fragmented message. Using the first fragment, the resolver deduces the size of the actual response and makes explicit requests (e.g., in parallel) for all other fragments (which can be deduced from the first fragments). To indicate to the name server that a particular query is for a fragment of a previous query, the resolver makes subsequent queries as ‘?n?domain_name’, where ‘?’ is a delimiter (a non-valid character for a domain, indicating that query or response is a fragment), ‘n’ is the number of the fragment, and ‘domain_name’ is the domain name of the actual query. Once the ‘n’ received by the resolver gets equal to the total number of fragments deduced by the resolver from the first fragment, all the fragments are re-assembled into the complete response and then verified.
@@
The issue with both aforementioned approaches is that none of them can accommodate double signatures as envisioned in this work for the security benefits discussed in Section 3. To accommodate two signatures while performing fragmentation at the application layer, herein, we devise an approach akin to QBF since it adopts standard RRs as discussed above. Fragmentation with double signatures poses a challenge because using a combination of pre- and post-quantum signatures and public keys results in ’A’ and ’DNSKEY’ responses that make it difficult for the name servers to create a first fragment with sufficient information. This first fragment is needed by resolvers to determine how many more fragments to expect before the response can be re-assembled and verified.
As per our analysis, the typical responses involved in address resolution with double signatures look something like those shown in Figure 8. Since the sizes of pre-quantum signatures and public keys (e.g., for ECDSA256, RSASHA256) are much shorter than post-quantum candidates (i.e., FALCON, DILITHIUM, SPHINCS+) it is inefficient to fragment both types of RRSIGs and DNSKEYs included in typical DNSSEC responses. Therefore, we ensure that entire pre-quantum RRSIGs and DNSKEYs are always sent in the first fragment. However, our observation is that this approach works well when we are dealing with the responses to ‘QTYPE’ of ‘A’, leaving sufficient space in the first fragment to include a fraction of post-quantum RRSIGs. The inclusion of the entire pre-quantum and a fraction of post-quantum RRSIG in the first fragment helps the resolver in estimating the number of subsequent fragments and making explicit requests for them in ‘parallel’ for efficiency reasons. However, our thorough analysis of the sizes of different response messages that typically make part of the address resolution process suggests that this fragmentation approach does not bode well with DNSKEY responses in certain instances (such as when RSASHA256 is used as a pre-quantum signature in double-signed DNSSEC). For example, the size of each field of all sections of a DNSKEY response to a query for ‘test.example’ is shown in Figure 9. It can be observed that the total size of this response is 4462B. Note that, this response is when RSASHA256 is combined with FALCON512 - i.e., the algorithm with the smallest post-quantum signatures. The other combinations will lead to even larger response sizes. In this (and other similar) case(s), sending the pre-quantum RRSIGs and DNSKEYs along with the header, question section, and additional section (see marked fields in Figure 9) in the first fragment takes up around 1178B. This leaves only 54B for sending any post-quantum RRSIGs and DNSKEYs, which is insufficient. In an instance like this, it becomes entirely difficult for a resolver to estimate the remaining number of fragments and thus can not make explicit requests for them.
@@
To circumvent the aforementioned issue, we leverage z-bits (which are reserved as of yet) in the header of a DNSSEC response. Whenever a name server is unable to send any part of post-quantum signatures or keys in the first fragment (e.g., with RSASHA256), it sets z-bits in the header such that they represent one out of three post-quantum signature algorithms (i.e., FALCON, DILITHIUM, SPHINCS+). Upon receiving a response, the resolver first checks whether z-bits are set or not. If z-bits are set, it first deduces the post-quantum algorithm and then by leveraging pre-quantum RRs that it has already received, estimates the total size of the response and the number of fragments this response should be divided into. Subsequently, the resolver makes parallel requests explicitly for all remaining fragments.
@@
Once all the fragments are received, the resolver first creates an updated message with both pre- and post-quantum RRSIGs and DNSKEYs and then combines all other fragments into this updated message. Our thorough analysis suggests that the aforementioned approach works accurately when various combinations of double signatures are incorporated in DNSSEC. However, we noticed that the re-assembled messages were not verified on the resolver (based on OQS-BIND). Our in-depth investigation suggested that this is due to the reason that DNS(SEC) uses a compression scheme for reducing the size of messages by eliminating the repetition of domain names in response messages. This is accomplished by replacing the domain names with a pointer to a previous occurrence of the same domain name (see RFC 1035 (Mockapetris, 1987)). Due to this, our observation is that when we always send pre-quantum RRSIGs and DNSKEYs in the first fragment, it sometimes changes the order of RRs in the reassembled message (as compared with the original message created by the name server). Therefore, this issue leads to a bad compression pointer due to which the responses fail to verify on the resolver.
@@
As a way of dealing with this issue, we use an offset to TTLs of RRs (see Fig. 3) in DNSKEY responses to keep track of the position of pre-quantum RRs in the original response messages on name servers. Once all the fragments are received by the resolver, it leverages this offset to create an updated message that places all the RRs in their original position. Our analysis suggests that this offset approach accurately re-assembles the messages such that they are correctly verified on the resolver.
@@
The overall working of double-signed fragmented DNSSEC is shown in Figure 10. As discussed above, whenever the response size is greater than the advertised threshold, the name server first estimates how many fragments this particular response should be divided into, sets the z-bits if needed and includes an offset to TTL of relevant RRs, and finally sets TC flag to ‘1’ to indicate to the resolver that this message is fragmented. The nameserver then sends the first fragment to the resolver and caches all other fragments. When the resolver receives this first fragment, it extracts the relevant information from this fragment to estimate how many additional fragments it should be expecting before the original message can be re-assembled. Once the fragments are estimated, the resolver then creates a separate query for all remaining fragments and sends these queries in parallel for efficiency reasons. For example, in Figure 10, we have assumed that the original response was divided into four fragments and then the resolver makes explicit queries in parallel for fragments 2 - 4 (i.e., ?2?test.socrates - ?4?test.socrates). Note that these fragments are formatted in the same way as in (Rawat and Jhanwar, 2024). When the name server receives these fragment requests, it sends the fragments that it has previously cached to the resolver. Once the resolver receives all the fragments (i.e., 4 in this representative example), it starts creating an updated message by utilizing TTL offset and re-assembles all the fragments. Finally, it re-sets TC and z-bits to ‘0’ and then attempts to verify the signatures to ensure the authenticity and integrity of the received IP addresses.

### v2

@@ 3.2. Quantum-Safe DNSSEC
Simply replacing the conventional digital signatures with post-quantum candidates mentioned above may not make the DNSSEC quantum-safe and may even make DNSSEC susceptible to attacks conducted using today’s classical computer (see the details mentioned above). Therefore, according to ENISA in (ENISA, 2022), a better and more cautious way to proceed is to use post-quantum digital signatures as an extra layer of protection in conjunction with conventional digital signatures. The combination essentially leads to a double-signed DNSSEC (note that the terms double-signed, combiners, or hybrid cryptosystems are used interchangeably). A well-designed double-signed DNSSEC will offer the benefit that it will remain strong if at least one of the used digital signatures is secure against a specific attack. In a double-signed DNSSEC, we will accept the IP address in a resolution process only if both signatures are verified. In particular, any attack revealed on post-quantum digital signatures in the near future that can be conducted with today’s classic computer will almost always fail to compromise the address resolution process, as the security of conventional digital signatures is established and assured against classical computers. Similarly, if the post-quantum digital signatures remain secure and no attacks can be conducted upon them, they will be secure against CRQC-driven attacks. Therefore, the double-signed approach is best suited for the interim period, as otherwise (i.e., with conventional or post-quantum digital signatures alone), DNSSEC may unknowingly be susceptible to attacks in the interim period as mentioned above.
@@
As far as the double signing is concerned, existing literature specifies two common interfaces (ENISA, 2022). The first one, referred to as detached-message interface, simply concatenates a pre-quantum signature $X$ with a post-quantum signature $Y$. The verifier only accepts the results when the $X$ verifier accepts the $X$ signature and the $Y$ verifier accepts the $Y$ signature. The second interface, referred to as signed-message interface, typically signs with a pre-quantum algorithm to get the first signature $X$ and then signs the result - i.e., $X$ with the post-quantum algorithm to get the final signature (e.g, $Z$). The double-signature verifier in this interface normally applies the post-quantum verifier first, followed by the pre-quantum verifier (or vice versa, depending upon the signing agreement).
@@
For the double-signed DNSSEC, we in this paper adopt an approach akin to the detached-message interface. Our reason is that this approach is relatively easier than others to integrate into our BIND9-based setup (see details on evaluation setup in Section 5). The overview of our double-signed DNSSEC is shown in Figure. 7. It essentially integrates two signatures - i.e., pre-quantum and post-quantum in the signed zone files on the authoritative server. Both these signatures are verified independently on the resolver before accepting any received resource record as verified, and trusted to offer enhanced security.
@@ 4. Double-Signatures and Maximum Message Size Issue
In this section, we discuss the message size issue when double signatures are incorporated in DNSSEC, along with the fragmentation approach that we have devised and tested for accommodating the dual signatures in DNSSEC.
@@
Since DNS(SEC) is an application-layer protocol, the generated messages are first placed in UDP packets, which subsequently are placed in IP packets that make the payload of link-layer frames. The size of these payloads has a limit that is dictated by the Maximum Transmission Unit (MTU) of the physical links that these frames traverse. The size of the frame’s payload is greater than the link’s MTU generally enforces routers to fragment these frames into smaller IP packets that travel separately to the resolver. However, the underlying transport protocol - i.e., UDP is mainly unreliable for the successful delivery of all these fragments and thus may lead to failure of address resolution. To circumvent this issue, either we need to resort to a reliable transport protocol - i.e., TCP, or we need to formulate messages such that their size is less than the MTU of the majority of physical links over which these DNS(SEC) messages will likely travel. Since, as per our discussion in Section 1, TCP has support issues in the existing DNS infrastructure and also is not as lightweight as UDP (Van Den Broek et al., 2014; van Rijswijk-Deij et al., 2016; Mao et al., 2022), the only option is to restrict DNS message size such that IP-level fragmentation is not triggered. To accomplish this, initial RFCs on DNS (Mockapetris, 1987) specify the maximum DNS message size to be less than (or equal to) 512B. However, the Extension Mechanism for DNS - i.e., EDNS(0) as defined in RFC 6891 (Damas and Vixie, 2013), theoretically allows a maximum message size of up to 64KB. Abstractly, this mechanism introduces a pseudo-RR called OPTION (alluded as OPT) and allows resolvers to specify buffer size in the ‘Class’ field of this pseudo-RR in the ‘additional section’ of an RR (see RR format in Fig. 3). Therefore, through EDNS(0), it is theoretically possible to increase DNS message size to a maximum of XB if the MTU of the end-to-end path between resolver and name server is at least ‘X+8(UDP Header)+40(IPv6 Header). However, there is no standard way to estimate the MTU of end-to-end links. Therefore, taking the cautious approach, the maximum DNS message size is limited to 1232B so as to avoid IP-level fragmentation and frequent TCP re-transmission (Manual, 2023) (assuming MTU of 1280B).
@@
The above-mentioned limit on the maximum DNS message size (i.e., 1232B) creates a particular challenge for double-signed DNSSEC as envisioned in this work. Precisely, any combination of pre-quantum and post-quantum signatures in double-signed DNSSEC will lead to response messages of size significantly larger than the limit discussed above. For example, our investigation shows that the message in response to a query of ‘Type A’ (see address resolution process depicted in Figure 7) is generally the smallest in size. However, our thorough analysis suggests that double-signed DNSSEC with pre- and post-quantum algorithms that lead to the signatures of the smallest sizes (e.g., pre-quantum ECDSA of 64B and post-quantum FALCON512 of 690B), has an approximate response to this query of size 2500B (note that exact size depends upon variable fields in an RR). The other messages, such as those that carry ‘DNSKEY RRs’, are even larger. Therefore, to accommodate double signatures in DNSSEC, we need to fragment these large responses at the application layer - i.e., within the DNS(SEC) protocol itself.
@@
There exist ways to fragment DNSSEC messages at the application layer in the post-quantum era. For example, ARRF (A Resource Record Fragment) mechanism (Goertzen and Stebila, 2022) proposes a request-based fragmentation approach at the application layer. For response messages larger than the threshold limit, ARRF on name servers ‘truncate’ the first fragments, indicating to the resolver that there are additional fragments before this response can be re-assembled and considered complete. To facilitate the resolver in efficient reception of fragments, ARRF introduces a new pseudo-RR called Resource Record Fragments (RRFRAGS, akin to OPT described above). This pseudo-RR adopts the format of a typical RR as shown in Figure 3; however, its fields are customized such that the name server can indicate to the resolver in the first fragment how the complete response would look like once all the fragments are received. Likewise, it also facilitates name servers in ascertaining whether the resolver requests a fragment (or is it a normal query). Although ARRF seems to perform fragmentation reliably, it has two main shortcomings. The first one is that it relies on a non-standard RR that will potentially cause issues with the middleboxes. Additionally, as acknowledged by the authors of ARRF, this approach may also be susceptible to memory exhaustion attacks.
@@
Given the above-mentioned problems with ARRF, QName-Based Fragmentation (QBF) is proposed in (Rawat and Jhanwar, 2024). Like ARRF, QBF also adopts a request-based fragmentation approach but leverages standard RRs, unlike ARRF, which relies on customized RRs for its operations. When the response size is bigger than the threshold, QBF sends a fraction of the signature or public key in the first fragment and sets the ‘TC’ flag in the header of the response to ‘1’, indicating to the resolver that this response is part of a fragmented message. Using the first fragment, the resolver deduces the size of the actual response and makes explicit requests (e.g., in parallel) for all other fragments (which can be deduced from the first fragment). To indicate to the name server that a particular query is for a fragment of a previous query, the resolver makes subsequent queries as ‘?n?domain_name’, where ‘?’ is a delimiter (a non-valid character for a domain, indicating that query or response is a fragment), ‘n’ is the number of the fragment, and ‘domain_name’ is the domain name of the actual query. Once the ‘n’ received by the resolver gets equal to the total number of fragments deduced by the resolver from the first fragment, all the fragments are re-assembled into the complete response and then verified.
@@
The issue with both aforementioned approaches is that none of them can accommodate double signatures as envisioned in this work for the security benefits discussed in Section 3. To accommodate two signatures while performing fragmentation at the application layer, herein, we devise an approach akin to QBF since it adopts standard RRs as discussed above. Fragmentation with double signatures poses a challenge because using a combination of pre- and post-quantum signatures and public keys results in ’A’ and ’DNSKEY’ responses that make it difficult for the name servers to create a first fragment with sufficient information. This first fragment is needed by resolvers to determine how many more fragments to expect before the response can be reassembled and verified.
As per our analysis, the typical responses involved in address resolution with double signatures look something like those shown in Figure 8. Since the sizes of pre-quantum signatures and public keys (e.g., for ECDSA256, RSASHA256) are much shorter than post-quantum candidates (i.e., FALCON, DILITHIUM, SPHINCS+), it is inefficient to fragment both types of RRSIGs and DNSKEYs included in typical DNSSEC responses. Therefore, we ensure that entire pre-quantum RRSIGs and DNSKEYs are always sent in the first fragment. However, our observation is that this approach works well when we are dealing with the responses to ‘QTYPE’ of ‘A’, leaving sufficient space in the first fragment to include a fraction of post-quantum RRSIGs. The inclusion of the entire pre-quantum and a fraction of post-quantum RRSIG in the first fragment helps the resolver in estimating the number of subsequent fragments and making explicit requests for them in ‘parallel’ for efficiency reasons. However, our thorough analysis of the sizes of different response messages that typically make part of the address resolution process suggests that this fragmentation approach does not bode well with DNSKEY responses in certain instances (such as when RSASHA256 is used as a pre-quantum signature in double-signed DNSSEC). For example, the size of each field of all sections of a DNSKEY response to a query for ‘test.example’ is shown in Figure 9. It can be observed that the total size of this response is 4462B. Note that, this response is when RSASHA256 is combined with FALCON512 - i.e., the algorithm with the smallest post-quantum signatures. The other combinations will lead to even larger response sizes. In this (and other similar) case(s), sending the pre-quantum RRSIGs and DNSKEYs along with the header, question section, and additional section (see marked fields in Figure 9) in the first fragment takes up around 1178B. This leaves only 54B for sending any post-quantum RRSIGs and DNSKEYs, which is insufficient. In an instance like this, it becomes entirely difficult for a resolver to estimate the remaining number of fragments and thus can not make explicit requests for them.
@@
To circumvent the aforementioned issue, we utilize z-bits (as specified in the original RFC 1035) in the header of a DNSSEC response. Note that since two of three z-bits are used in RFC2535, however, since we adopt a daemon to accurately fragment and reassemble packets, we can still use these bits for fragmentation and reassembly in the daemon (see details of our implementation in Section 5). Whenever a name server is unable to send any part of post-quantum signatures or keys in the first fragment (e.g., with RSASHA256), it (i.e., daemon) sets z-bits in the header (of the first fragment) such that they represent one out of three post-quantum signature algorithms (i.e., FALCON, DILITHIUM, SPHINCS+). Upon receiving a response (i.e., first fragment), the resolver daemon first checks whether z-bits are set or not. If z-bits are set, it first deduces the post-quantum algorithm and then, by leveraging pre-quantum RRs that it has already received, estimates the total size of the response and the number of fragments this response should be divided into. Subsequently, the resolver makes parallel requests explicitly for all remaining fragments.
@@
Once all the fragments are received, the resolver first creates an updated message with both pre- and post-quantum RRSIGs and DNSKEYs and then combines all other fragments into this updated message with header values exactly as in the original message (which can be taken from the last fragment, as in the first fragment, we repurposed the z-bits). Our thorough analysis suggests that the aforementioned approach works accurately when various combinations of double signatures are incorporated in DNSSEC. However, we noticed that the re-assembled messages were not verified on the resolver (based on OQS-BIND). Our in-depth investigation suggested that this is because DNS(SEC) uses a compression scheme to reduce the size of messages by eliminating the repetition of domain names in response messages. This is accomplished by replacing the domain names with a pointer to a previous occurrence of the same domain name (see RFC 1035 (Mockapetris, 1987)). Due to this, our observation is that when we always send pre-quantum RRSIGs and DNSKEYs in the first fragment, it sometimes changes the order of RRs in the reassembled message (as compared with the original message created by the name server). Therefore, this issue leads to a bad compression pointer, due to which the responses fail to verify on the resolver.
@@
As a way of dealing with this issue, we use an offset to TTLs of RRs (see Fig. 3) in DNSKEY responses to keep track of the position of pre-quantum RRs in the original response messages on name servers. Once all the fragments are received by the resolver, it leverages this offset to create an updated message that places all the RRs in their original position. Our analysis suggests that this offset approach accurately reassembles the messages such that they are correctly verified on the resolver.
@@
The overall working of double-signed fragmented DNSSEC is shown in Figure 10. As discussed above, whenever the response size is greater than the advertised threshold, the daemon on name server first estimates how many fragments this particular response should be divided into, sets the z-bits if needed, and includes an offset to TTL of relevant RRs, and finally sets the TC flag to ‘1’ to indicate to the resolver that this message is fragmented. The nameserver then sends the first fragment to the resolver and caches all other fragments. When the resolver receives this first fragment, it extracts the relevant information from this fragment to estimate how many additional fragments it should expect before the original message can be reassembled. Once the fragments are estimated, the resolver then creates a separate query for all remaining fragments and sends these queries in parallel for efficiency reasons. For example, in Figure 10, we have assumed that the original response was divided into four fragments and then the resolver makes explicit queries in parallel for fragments 2 - 4 (i.e., ?2?test.socrates - ?4?test.socrates). Note that these fragments are formatted in the same way as in (Rawat and Jhanwar, 2024). When the name server receives these fragment requests, it sends the fragments that it has previously cached to the resolver. Once the resolver receives all the fragments (i.e., 4 in this representative example), it starts creating an updated message by utilizing TTL offset and reassembles all the fragments. Finally, it resets TC and makes the other header values as in the original message (which can be copied from the last fragments). Once done, the resolver then attempts to verify the signatures to ensure the authenticity and integrity of the received IP addresses.